OpenStudy (anonymous):

one canned juice drink in 20 % orange juice; another is 10 % orange juice. How many liters of each should be mixed together in order to get 10 L thats is 18% orange juice ? please and thank you !!:))

OpenStudy (anonymous):

(20/100)*x + (10/100)*(10-x) = (18/100)*10 2x/10 + 1 - x/10 = 1.8 x/10 = 0.8 x=8 so 8L of the 20% and 2L of the 10%